The length of a rectangle is three times its width. The area of the rectangle is 170 square meters. Find the width?
Answer by checkley79(3341)

L=3W
AREA=LW
170=3W*W
170=3W^2
W^2=170/3
W^2=56.67
W=SQRT56.67
W=7.53 METERS FOR THE WIDTH.
L=3*7.53
L=22.58 ANS FOR THE LENGTH.
PROOF:
170=22.58*7.53
170=170